# Code Generation with Small Language Models: A Deep Evaluation on Codeforces

### v1

@@ III-D Consistency Analysis
While pass@k measures how often a model produces at least one correct solution, it does not capture consistency across multiple attempts. To address this, we use Semantic Consistency (SC), which evaluates how reliably a model reproduces correct outputs in repeated submissions [29]. This metric is particularly important for code generation, where consistent correctness matters more than isolated success. As shown in Table I, all models achieved SC above 60%, with Gemma 2 9B and Gemma 3 12B leading despite lower overall accuracy. Phi-4 14B also showed strong consistency (77.5%), indicating that once a model solves a problem, it tends to do so repeatedly—a sign of learned, stable behavior rather than random chance.
III-E Phi-4 14B: C++
@@
To investigate whether the choice of programming language influences model performance, we evaluated Phi-4 14B, which has the best performance in our evaluation, using both Python and C++. C++ is widely adopted in competitive programming due to its performance efficiency, particularly in time-critical problems. In this experiment, we kept the prompting strategy, modifying only the target language.
@@
The results show that Phi-4 14B achieves comparable accuracy in both languages: 63.6% pass@3 in Python and 61.1% in C++. However, the sets of problems solved in each language differ around 20%. The Python version solved 35 problems not solved in C++, while the C++ version succeeded on 28 problems not handled in Python. When combining both sets (see Figure 2), Phi-4 14B solves 206 out of 280 problems—achieving an aggregated pass@1 of 63.6%, pass@2 of 70.4%, and pass@3 of 73.6%.
@@
Of the 28 problems uniquely solved in C++, 25 are rated above 1500 and primarily fall under categories such as Dynamic Programming, Math, Greedy, Brute Force, and Graphs. In these cases, Python submissions often failed due to Time Limit Exceeded (12), Wrong Answers (18), or Runtime Errors (9). These findings suggest that problem complexity and computational constraints may favor C++ over Python in certain scenarios.
[…]
III-F Larger Models
@@
To understand how far SLMs are from the best-performing models, we analyzed OpenAI’s proprietary o3-mini-high, which, as shown in Table I, achieved the highest overall accuracy with a pass@3 of 86.8%. Phi-4 14B successfully solved five problems that were not answered by o3-mini-high. These problems had Elo ratings of 1400, 1500, 1900 (2x) and 2100. To further explore its limitations, we submitted the 38 problems that o3-mini-high failed to solve to OpenAI’s larger reasoning model, o1 [7], using three attempts per problem. These problems ranged from Elo 1400 to 2100 and focused on challenging categories such as Dynamic Programming, Math, Trees, and Data Structures. o1 successfully solved only 3 of the 38 problems (7.9% pass@3). Among the three successful cases, problem 479-C (Elo 1400) was solved by both o1 and Phi-4 14B, but not by o3-mini-high. The other two solutions involved complex Data Structures at Elo 2100. For the remaining 35 problems, the failures were mainly due to Time Limit Exceeded (30 cases), Wrong Answers (14), Memory Limit Exceeded (5), and one Runtime Error. These findings confirm that even state-of-the-art proprietary models struggle with high-difficulty competitive programming tasks.
@@
However, when aggregating Phi-4 14B’s results across Python and C++ (Section III-E), its pass@3 increases to 73.6%, narrowing the gap to o3-mini-high. This highlights the potential of open models to achieve competitive performance through multi-language strategies. Both models follow a similar trend: strong accuracy on easier problems with a gradual decline as difficulty increases (Figure 2). Moreover, Phi-4 14B is open, cost-effective, and runs on consumer-grade hardware, making it an attractive alternative for a wide range of applications.
[…]
III-G Qualitative Analysis
@@
To complement our quantitative evaluation metrics such as pass@k and semantic consistency, we conducted a qualitative analysis of the errors made by Phi-4 14B. One of the authors, a competitive programmer with over seven years of experience, has reached a peak ELO rating of 2245 on Codeforces. He conducted a manual review of a sample of 21 (~$\texttildelow 20\%$) problems that Phi-4 14B failed to solve. Based on this analysis, we identified four primary causes of failure: near-correct solutions with minor mistakes, algorithmically correct solutions that exceeded time limits, incorrect solutions with relevant direction, and fundamentally flawed implementations.
@@
The first category of failures involved near-correct solutions that required only minor adjustments to succeed. In five of the reviewed problems, Phi-4 14B produced almost correct code, needing simple fixes such as correcting variable initialization (Problem 1343-C), handling edge cases (Problem 1399-C), or adjusting state transitions in dynamic programming (Problem 698-A). Three additional problems required more modifications. In Problem 489-C, multiple assignment errors and output formatting issues had to be corrected. Problem 1-B involved a minor logical issue, which required adding a loop to properly parse the input. In Problem 161-D, the Python solution was logically sound but inefficient; performance improvements were necessary, including replacing dictionaries with lists and rewriting recursive functions as iterative ones to meet time constraints.
@@
In the second category, three problems featured conceptually correct algorithms that failed due to time constraints. When we prompted Phi-4 14B to rewrite these solutions in C++—preserving the logic but aiming for improved efficiency—two were successfully solved. However, Problem 1541-B (Elo 1200) remained unsolved, as it required a reduction in time complexity beyond what the model could infer.
[…]
@@
In 4 of the reviewed problems, Phi-4 14B produced outputs that were in the right direction but still far from being correct. These solutions often included algorithmic elements aligned with the correct approach—such as identifying the right data structure or outlining the high-level logic—but lacked crucial implementation details or contained major logical errors. While the model appeared to grasp the core idea behind the problems, its execution was flawed and incomplete. As a result, these attempts were not easily repairable and required substantial rewriting to become viable. Still, one noteworthy aspect is that this type of partially correct attempt can offer valuable insights to the user. In problems where it is particularly hard to identify a viable direction, having the model suggest an initial high-level approach—even if incomplete—can be quite helpful as a starting point for further reasoning. In Problem 339-D (Elo 1700), the failure was attributed to incomplete problem description caused by formatting issues during automatic extraction. For example, numbers with exponents were incorrectly ignored, altering the meaning of the problem statement. As future work, we plan to enhance the text extraction process to avoid such issues.
@@
Finally, in 5 of the 21 cases, the model produced fundamentally incorrect code. These outputs were either logically incoherent or based on an incorrect understanding of the problem statement. In some cases, the specialist reviewer noted that “the SLM seems not to have understood the problem at all.” In all three attempts, the model made incorrect assumptions and attempted to solve a completely different task.
[…]
IV Conclusion
@@
This study evaluated the performance of five open SLMs—Llama 3.2 3B, Gemma 2 9B, Gemma 3 12B, DeepSeek-R1 14B, and Phi-4 14B—on 280 competitive programming problems from Codeforces. Phi-4 14B stood out with a pass@3 of 63.6% and semantic consistency of 77.5%, significantly outperforming other SLMs. In contrast, smaller models like Llama 3.2 3B and Gemma 2 9B achieved pass@3 rates below 24%. We also evaluated Phi-4 14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@3 to 73.6%. For comparison, the proprietary o3-mini-high model achieved a pass@3 of 86.8%.
@@
A manual review of incorrect Phi-4 14B submissions revealed that most solutions were close to correct, with errors arising from minor issues—such as unhandled edge cases or suboptimal implementations—rather than fundamental reasoning flaws. Notably, 26 problems that failed in Python were successfully solved in C++, suggesting that leveraging multiple programming languages can enhance model performance. These results show that SLMs—particularly Phi-4 14B—strike a compelling balance between performance, efficiency, and accessibility. They represent a practical alternative to large proprietary models, especially in settings with limited resources.
While previous works have analyzed the performance of models like AlphaCode [9] and o3-mini-high [7] on Codeforces, focusing primarily on identifying correct and incorrect answers in LLMs, our study goes further by examining the performance of SLMs across various topics, Elo ratings, and providing a detailed qualitative analysis of the errors. Future work may explore higher-ELO problems, refined prompting strategies, alternative configurations, language-specific generation, and lightweight tuning techniques to further assess and improve model effectiveness.

### v2

@@ III-E Consistency Analysis
While pass@k measures how often a model produces at least one correct solution, it does not capture consistency across multiple attempts. To address this, we introduce Semantic Consistency (SC), which evaluates how reliably a model reproduces correct outputs in repeated submissions [25]. This metric is particularly important for code generation, where consistent correctness matters more than isolated success. As shown in Table I, all models achieved SC above 60%, with Gemma-3-12B leading despite lower overall accuracy. Phi-4-14B also showed strong consistency (77.5%), indicating that once the model solves a problem, it tends to do so repeatedly—a sign of learned, stable behavior rather than random chance.
III-F Phi-4-14B: C++
@@
To investigate whether the choice of programming language influences model performance, we evaluated Phi-4-14B, which has the best performance in our evaluation, using both Python and C++. C++ is widely adopted in competitive programming due to its performance efficiency, particularly in time-critical problems. In this experiment, we kept the prompting strategy, modifying only the target language. The combined results reported represent the union of problems solved by either language, simulating an ideal scenario where solutions in both languages can be leveraged to maximize coverage.
@@
The results show that Phi-4-14B achieves comparable accuracy in both languages: 63.6% pass@3 in Python and 61.1% in C++. However, the sets of problems solved in each language differ around 20%. The Python version solved 35 problems not solved in C++, while the C++ version succeeded on 28 problems not handled in Python. The combined results—union of problems solved in either language—represent an upper-bound scenario where multi-language solutions maximize coverage (see Figure 2) In this case, Phi-4-14B solves 206 out of 280 problems—achieving an aggregated pass@2 of 63.6%, pass@4 of 70.4%, and pass@6 of 73.6%.
@@
Of the 28 problems uniquely solved in C++, 25 are rated above 1500 and primarily fall under categories such as Dynamic Programming, Math, Greedy, Brute Force, and Graphs. Python failures on these problems were often due to Time Limit Exceeded (12), Wrong Answers (18), or Runtime Errors (9), suggesting that differences are driven by computational constraints and problem complexity rather than the language itself.
Although developers may not always have the freedom to choose the programming language for a given task—due to platform restrictions, legacy systems, or deployment environments—multi-language generation remains a valuable strategy in several contexts. For instance, in latency-critical scenarios such as AWS Lambda, developers may prefer deploying solutions in more performant languages like C++ to reduce execution time. Recent work such as InterTrans [9] also highlights how leveraging multiple target languages can improve coverage and effectiveness in code generation pipelines, reinforcing the practical relevance of this approach.
III-G Larger Models
@@
To understand how far SLMs are from the best-performing models, we analyzed o3-mini-high, which, as shown in Table I, achieved the highest overall accuracy with a pass@3 of 86.8% on the same dataset. Phi-4-14B successfully solved five problems that were not answered by o3-mini-high. These problems had Elo ratings of 1400, 1500, 1900 (2x) and 2100. These findings confirm that even state-of-the-art proprietary models struggle with high-difficulty competitive programming tasks.
@@
However, when aggregating Phi-4-14B’s results across Python and C++ (Section III-F), its pass@3 increases to 73.6%, narrowing the gap to o3-mini-high. This highlights the potential of open models to achieve competitive performance through multi-language strategies. Both models follow a similar trend: strong accuracy on easier problems with a gradual decline as difficulty increases (Figure 2). Moreover, Phi-4-14B is open, cost-effective, and runs on consumer-grade hardware, making it an attractive alternative for a wide range of applications.
III-H Metamorphic Testing
To assess data contamination [4] in our benchmark, we conducted metamorphic testing on 14 problems with at least one submission accepted—one per difficulty level—using Phi-4-14B and Python language, the best-performing model. This process involved altering problem statements through minor edits such as renaming variables, replacing verbs, and using synonyms, without changing their core semantics. Before transformation, Phi-4-14B achieved pass@1/2/3 scores of 78.57%, 92.86%, and 100%. After transformation, the scores were 69.05%, 80.95%, and 85.71%, respectively. Besides the small drop in performance, these values still suggest no evidence of data contamination.
III-I Qualitative Analysis
@@
We conducted a qualitative analysis of the errors made by Phi-4-14B. One of the authors, a competitive programmer with over seven years of experience, has reached a peak ELO rating of 2245 on Codeforces. He conducted a manual review of a sample of 21 ($\texttildelow 20\%$) problems that Phi-4-14B failed to solve. Based on this analysis, we identified four primary causes of failure: near-correct solutions with minor mistakes, algorithmically correct solutions that exceeded time limits, incorrect solutions with relevant direction, and fundamentally flawed implementations.
@@
The first category of failures involved near-correct solutions that required only minor adjustments to succeed. In five of the reviewed problems, Phi-4-14B produced almost correct code, needing simple fixes such as correcting variable initialization (Problem 1343-C), handling edge cases (Problem 1399-C), or adjusting state transitions in dynamic programming (Problem 698-A). Three additional problems required more modifications. In Problem 489-C, multiple assignment errors and output formatting issues had to be corrected. Problem 1-B involved a minor logical issue, which required adding a loop to properly parse the input. In Problem 161-D, the Python solution was logically sound but inefficient; performance improvements were necessary, including replacing dictionaries with lists and rewriting recursive functions as iterative ones to meet time constraints.
@@
In the second category, three problems featured conceptually correct algorithms that failed due to time constraints. When we prompted Phi-4-14B to rewrite these solutions in C++—preserving the logic but aiming for improved efficiency—two were successfully solved. However, Problem 1541-B (Elo 1200) remained unsolved, as it required a reduction in time complexity beyond what the model could infer.
Many of Phi-4-14B’s incorrect answers were due to minor implementation issues or edge-case handling, rather than fundamental misunderstandings. These errors were often correctable with small edits. Addressing such cases systematically could bring the model’s performance even closer to that of proprietary models like o3-mini-high, reinforcing the value of open solutions in competitive programming tasks.
@@
In 4 of the reviewed problems, Phi-4-14B produced outputs that were in the right direction but still far from being correct. These solutions often included algorithmic elements aligned with the correct approach—such as identifying the right data structure or outlining the high-level logic—but lacked crucial implementation details or contained major logical errors. While the model appeared to grasp the core idea behind the problems, its execution was flawed and incomplete. As a result, these attempts were not easily repairable and required substantial rewriting to become viable. Still, one noteworthy aspect is that this type of partially correct attempt can offer valuable insights to the user. In problems where it is particularly hard to identify a viable direction, having the model suggest an initial high-level approach—even if incomplete—can be quite helpful as a starting point for further reasoning. In Problem 339-D (Elo 1700), the failure was attributed to incomplete problem description caused by formatting issues during automatic extraction. For example, numbers with exponents were incorrectly ignored, altering the meaning of the problem statement. As future work, we plan to enhance the text extraction process to avoid such issues.
@@
Finally, in 5 of the 21 cases, the model produced fundamentally incorrect code. These outputs were either logically incoherent or based on an incorrect understanding of the problem statement. In some cases, the specialist reviewer noted that “the model seems not to have understood the problem at all.” In all 3 attempts, the model made incorrect assumptions and attempted to solve a completely different task. To reduce costs, prioritize open models—Phi-4-14B in Python, then C++—and check outputs for minor implementation errors before using expensive proprietary options like o3-mini-high.
III-J Threats to Validity
This study is limited by the use of fixed prompts, quantization settings, models, and a dataset of 280 popular Codeforces problems, which may introduce topic and difficulty biases. Results may vary with alternative configurations or larger benchmarks. Manual error analysis by a single specialist may introduce bias and pass@k only captures success rates, not code quality aspects.
IV Related Work
Hou and Ji [3] evaluated LLMs, including GPT-4, Claude 2, LLaMA 2, GPT-3.5, Gemini Ultra, Gemini Pro, and Code LLaMA, on LeetCode and GeeksforGeeks challenges across varying difficulty levels, allowing up to five attempts per problem. Sakib et al. [6] focused on GPT-3.5 with stricter constraints (two attempts) across three difficulty levels and diverse algorithmic topics. Their evaluation covered three difficulty levels and a diverse range of algorithmic topics, offering insights into how well LLMs handle different problem-solving paradigms.
Jain et al.[5] proposed LiveCodeBench, a holistic, contamination-free evaluation framework for LLMs using Codeforces problems. Continuously updated with new problems from Codeforces and other platforms, it assesses a wide range of code-related capabilities, including code generation, self-repair, execution, and test case prediction. The study evaluated 18 base and 34 instruction-tuned LLMs, uncovering insights on dataset contamination, performance comparisons, and overfitting. Complementing benchmark-based evaluations, recent agentic-driven approaches such as self-collaboration[8] and self-planning [2] improve problem-solving by structuring tasks into subtasks or plans, achieving up to 47.1% gains over single-agent baselines. Large-scale analyses on platforms like Codeforces [24] further reveal model capabilities across topics and difficulties, e.g., GPT-Neo solving around 20% of introductory problems, highlighting incremental advances in code correctness.
While previous works have analyzed the performance of models like AlphaCode [18] and o3-mini-high[17] on Codeforces, focusing mainly on correctness, our study instead highlights the behavior of smaller models, covering multiple topics and Elo ratings and adding a qualitative error analysis. Recent benchmarks such as LiveCodeBench Pro [7] and Wei et al. [10] provide valuable insights into the limitations of large models across problem types and reasoning categories, and our work complements them by shifting the focus to SLMs with a more fine-grained evaluation.
V Conclusion
@@
This study evaluated the performance of three open SLMs—Llama-3.2-3B, Gemma-3-12B, and Phi-4-14B—on 280 competitive programming problems from Codeforces. Phi-4-14B stood out with a pass@3 of 63.6% and semantic consistency of 77.5%, significantly outperforming other SLMs. In contrast, smaller models like Llama-3.2-3B and Gemma-3-12B achieved pass@3 rates below 20%. We also evaluated Phi-4-14B on C++ code generation and found that combining outputs from both Python and C++ increased its aggregated pass@6 to 73.6%. For comparison, the proprietary o3-mini-high model achieved a pass@3 of 86.8%.
@@
A manual review of incorrect Phi-4-14B submissions revealed that most solutions were close to correct, with errors arising from minor issues—such as unhandled edge cases or suboptimal implementations—rather than fundamental reasoning flaws. Notably, 26 problems that failed in Python were successfully solved in C++, suggesting that leveraging multiple programming languages can enhance model performance. These results show that SLMs—particularly Phi-4-14B—strike a compelling balance between performance, efficiency, and accessibility. They represent a practical alternative to large proprietary models, especially in settings with limited resources. In environments with tighter computational constraints, smaller models such as Llama-3.2-3B, or Gemma-3-12B can still perform effectively on lower-difficulty problems (e.g., Codeforces ratings between 800 and 1000). Future work may explore higher-ELO problems, refined prompting strategies, alternative configurations, language-specific generation, and lightweight tuning techniques to further assess and improve model effectiveness.
As future work, we intend to include efficiency-oriented metrics such as inference time, latency, energy consumption, and cost-per-accepted solutions, which are critical for practical deployment but not assessed here. Contamination was only checked through metamorphic testing, and more rigorous approaches like time-split controls are needed. Broader baselines and ablation studies (e.g., temperature, seeds, quantization, self-repair) could also clarify configuration sensitivity, while automated pipelines for bug categorization would refine error analysis beyond the manual audit conducted in this work.
Acknowledgments
We thank the anonymous reviewers for their valuable feedback. This work was partially supported by CNPq (403719/2024-0, 310313/2022-8, 404825/2023-0, 443393/2023-0, 312195/2021-4), FAPESQ-PB (268/2025).